A geometric sequence is defined using the following rule
f(0)=4 and f(n)=f(n-1).-2
Find the fourth term of the sequence.show how you arrived at your answer

Answer by stanbon(75887)

A geometric sequence is defined using the following rule
f(0)=4 and f(n)=f(n-1).-2
Find the fourth term of the sequence.show how you arrived at your answer
-----
f(0) = 4
f(1) = 4.-2 (not sure what you mean by .-2)
f(2) = 4.-2*(.-2)
f(3) = 4.-2*(.-2)(.-2)
----------------
Cheers,
Stan H.